# Towards Trustworthy Vital Sign Forecasting: Leveraging Uncertainty for Prediction Intervals

### v1

@@ I Introduction
    Richer uncertainty-width mapping. We allow more expressive functions than simple scalar multiplication to capture complex relationships between uncertainty estimates $u(\mathbf{x})$ and interval half-width $\epsilon$.
@@ Appendix B Implementation Details
To further expedite the search process in step 1 of KNN PI, we employ a K-Dimensional Tree [56] fitted on the validation set reconstruction errors before inference. During inference, the tree is queried with the reconstruction errors of $\mathbf{x}$.
We applied RUE, MCD, and IN to an MLP prediction model and also used an MLP decoder for RUE. For RUE, we fine-tuned the width and depth of the encoder and decoder through grid search, using validation loss and correlation with loss as the performance metrics for the encoder and decoder, respectively. Regularisation on the hidden layers of the decoder and early stopping ($\text{patience}=20$) were implemented to mitigate overfitting. The hyperparameters of all other baselines were tuned with grid search in a similar manner, using validation performance as the metric. We repeated all experiments five times and reported the mean and standard deviation of all evaluation metrics.
For implementation details, hyperparameter settings, and ablation studies on the KNN parameter $k$, refer to the code repository: https://anonymous.4open.science/r/rue_prediction_interval-0DF4.
@@ Appendix C Evaluating Time Series Prediction Performance
Table II presents the average forecasting performance of various uncertainty estimation models across three time horizons. For the MIMIC dataset, SGPR achieved the best prediction performance, followed by IN and RUE (or generic multilayer perceptron (MLP)), as evidenced by their low MSE values across all three horizons. On the PhysioNet dataset, RUE, MCD, SGPR and IN had comparable performances. Across both datasets, BNN and DER exhibited the weakest performance. These results demonstrate that the choice of uncertainty estimation method can significantly influence prediction performance.
@@ Appendix D Evaluating Uncertainty Estimation Performance
We evaluate the uncertainty estimates across all three time horizons using the following three metrics. Note that in all three metrics, prediction loss is computed as the mean absolute error (MAE).
@@
The correlation between the uncertainty estimate and the prediction error [33]. A highly positive correlation suggests a reliable uncertainty estimate, as it indicates that when prediction errors are high, uncertainty estimates are also high, and vice versa. We also report the two-sided p-value of the correlation test to determine the significance of the observed correlation.
@@
The area under the risk-coverage curve [57], which plots prediction loss (Risk) against the proportion of confident predictions (Coverage) on the test set. A lower AURC indicates a more selective uncertainty estimate. Notably, AURC is invariant to changes in the prediction model, enabling comparisons of uncertainty estimates across models with varying predictive performance.
@@
This metric quantifies the prediction error of confident predictions (instances with normalized uncertainty estimates lower than $\sigma\in[0.1,0.2]$), measuring the robustness of the uncertainty estimate to false negatives (incorrect predictions with high confidence). A lower $\sigma$-risk score indicates greater resilience to false negatives.
To minimize the impact of outliers on the $\sigma$-risk score, we normalize the uncertainty estimates with the following procedure: (1) Compute the minimum and maximum uncertainty estimates after excluding instances with uncertainty values beyond 1.5 interquartile ranges from the third quartile; (2) Use these computed minimum and maximum values to apply min-max normalization to the uncertainty estimates.
TABLE III: Comparison of Uncertainty Estimation Performance on MIMIC and PhysioNet Across All Time Horizons. For each metric, we bold the values for the best-performing model and underline the values for the second-best-performing model.
@@
Table III compares the performance of each uncertainty estimate across both datasets and all time horizons. We observe that uncertainty estimation reliability generally decreases as the time horizon lengthens, indicated by the declining correlation between uncertainty estimates and prediction error from $t+1$ to $t+3$. This suggests that existing uncertainty estimates struggle to capture uncertainty effectively as the prediction horizon expands.
Comparing the correlation of each uncertainty estimate with prediction error, we find that RUE is the most reliable estimate by a significant margin on MIMIC, followed by BNN. On PhysioNet, RUE remains the most reliable, with DER as the next best estimate. The relative performance of RUE, BNN and DER suggests that distributional uncertainty is the predominant source of uncertainty in both datasets, whereas greater epistemic and aleatoric uncertainty is present in the MIMIC and PhysioNet, respectively.
Comparing AURC, we observe that SGPR is the most effective uncertainty estimate for selective prediction, followed closely by RUE, especially for time horizons beyond $t+1$. This demonstrates RUE’s potential to enhance prediction performance when used for selective prediction (where the model only makes predictions when it is confident), as RUE yields the lowest selective prediction loss (AURC) even when applied to models with lower overall performance (Table II).
Comparing $\sigma$-risk scores, we note that RUE is the most robust to false negatives, with the smallest or second smallest ($T+1$ for MIMIC) $\sigma$-risk score across most datasets and time horizons, followed by DER and SGPR. This suggests that confident predictions identified by RUE are less likely to be incorrect compared to other uncertainty estimates.

### v2

# Anonymous    1st Li Rong Wang College of Computing and Data Science

@@ I Introduction
    Richer uncertainty-width mapping. We allow more expressive nonlinear functions – such as Gaussian copulas and K-nearest neighbours (KNN) —- beyond simple scalar multiplication to capture complex relationships between uncertainty estimates $u(\mathbf{x})$ and interval half-width $\epsilon$.
@@ Appendix B Implementation Details
To expedite the search process in step 1 of KNN PI, we employ a K-Dimensional Tree [56] fitted on validation set reconstruction errors before inference, and queried with the reconstruction errors of $\mathbf{x}$ during inference. We applied RUE, MCD, and IN to an MLP prediction model, using an MLP decoder for RUE. For RUE, encoder and decoder width/depth were fine-tuned via grid search, with validation loss and correlation with loss as their respective metrics. Regularisation and early stopping ($\text{patience}=20$) mitigated overfitting. Hyperparameters of other baselines were tuned similarly via grid search using validation performance. All experiments were repeated five times, reporting mean and standard deviation of all metrics. For implementation details, hyperparameter settings, and ablation studies on the KNN parameter $k$, refer to the code repository: https://github.com/lr98769/rue_prediction_interval.
@@ Appendix C Evaluating Time Series Prediction Performance
Table II presents the average forecasting performance of various uncertainty estimation models across three time horizons. For the MIMIC dataset, SGPR achieved the best prediction performance, followed by IN and RUE (or multilayer perceptron (MLP)), as evidenced by their low MSE values across all horizons. On the PhysioNet dataset, RUE, MCD, SGPR and IN had comparable performances. Across both datasets, BNN and DER exhibited the weakest performance. These results demonstrate that the choice of uncertainty estimation method can significantly influence prediction performance.
@@ Appendix D Evaluating Uncertainty Estimation Performance
We evaluate uncertainty estimates using three metrics, with prediction loss measured by mean absolute error (MAE).
@@
The correlation between the uncertainty estimate and prediction error [33], where a strong positive correlation indicates a reliable estimate.
@@
The area under the risk-coverage curve [57], which plots prediction loss (risk) against the proportion of confident predictions (coverage). Lower AURC indicates a more selective uncertainty estimate.
@@
Measures prediction error for confident predictions (normalized uncertainty $\leq\sigma\in[0.1,0.2]$), reflecting robustness to false negatives (incorrect predictions with high confidence). Lower values indicate greater resilience. To mitigate the impacts of outliers, uncertainty is normalized by (1) excluding values beyond 1.5 IQR above Q3 and (2) applying min–max normalization on the remaining range.
TABLE III: Comparison of Uncertainty Estimation Performance on MIMIC and PhysioNet Across Time Horizons. For each metric, we bold values for the best-performing model and underline values for the second-best-performing model.
@@
Table III compares uncertainty estimate performance across datasets and horizons. Reliability generally decreases with longer horizons, as shown by the declining correlation with prediction error from $t+1$ to $t+3$, suggesting existing estimates struggle to capture uncertainty at extended horizons.
Correlation: RUE is the most reliable estimate on both MIMIC and PhysioNet, followed by BNN on MIMIC and DER on PhysioNet.
AURC: SGPR achieves the lowest AURC, followed closely by RUE, particularly beyond $t+1$. This highlights RUE’s potential for selective prediction (Table II).
$\sigma$-Risk Scores: RUE is the most robust to false negatives, showing the smallest or second-smallest $\sigma$-risk across datasets and horizons, followed by DER and SGPR, indicating its confident predictions are less likely to be incorrect.